OpenStudy (anonymous):

ACT math practice question. There are three colors of jelly beans in a bowl. If there are 5 more red ones than blue, and there are 9 more blue ones than green, and the total number of beans is 26, how many green ones are in the bowl? Now I know there should be some equation set up that I can then solve, I just can't think of what the equation would be for some reason. Please help!

OpenStudy (thomas5267):

Let the number of red balls be r, the number of blue balls be b and the number of green balls be g. Try set up a set of equations.

OpenStudy (anonymous):

All I got is b + 5 = r and b - 9 = g But I don't know how that is solvable. Not enough information.

OpenStudy (anonymous):

how many total jelly beans?

OpenStudy (anonymous):

26

OpenStudy (anonymous):

what would be the equation for that total?

OpenStudy (anonymous):

in terms of the variables...

OpenStudy (anonymous):

b + r + g = 26. But I have a feeling that's not it... Would (b + 5) + (b - 9) + b = 26 work? Even then I don't know how that is solvable either.

OpenStudy (anonymous):

yes, and yes... to solve the last eq you wrote, comine like terms, isolate the variable and voila!

OpenStudy (anonymous):

"combine"

OpenStudy (anonymous):

Oh! Okay, I think it is coming back to me. It's just been a while since I've done this. So to combine like terms, before simplifying it, would I get 2b - 9b + 5b -45 + b = 26 ??

OpenStudy (anonymous):

Oh wait, that can't be it because I just got that b = -26 ...

OpenStudy (anonymous):

not quite... (b+5)+b+(b-9) = 29 =>b+b+b+5-9=26 => 3b - 4 = 26 can you solve from there?

OpenStudy (anonymous):

26 not 29

OpenStudy (anonymous):

Okay, thanks I can!

OpenStudy (anonymous):

what do you get? for r, b and g?

OpenStudy (anonymous):

\[green=x\]\[blue=x+9\]\[red=(x+9)+5=x+14\]\[g+b+r=x+x+9+x+14=3x+23\]\[3x+23=26\]\[x=1\]

OpenStudy (anonymous):

\[green=x=1\]\[blue=x+9=1+9=10\]\[red=x+14=1+14=15\]\[1+10+15=26\] as given

OpenStudy (error1603):

r